CALL FOR PAPERS

8^{th} International Workshop on High Performance and Distributed Mining (HPDM’05)

in conjunction with the

5^{th} International SIAM Conference on Data Mining (SDM’05) April 23, 2005

Over the years the definition of high performance computing has taken on various forms as a function of the types of technical and creative uses and the underlying semantics of the applications driving them. Traditional definitions often refer to the problem of using high end parallel computers to meet the need of scientific applications. However, high performance computing can also include the need for fast sequential algorithms that target memory and I/O performance. The last decade has seen the growth and importance of grid computing where resources and data are physically distributed. This has led to the development of high performance distributed algorithms over the computational grid, where privacy, security, and resource discovery are all important issues. This year the workshop welcomes papers on all aspects of high performance data mining.

Topics of interest include (but are not limited to):

• Grid-based data mining algorithms and systems.

• Distributed techniques for incremental, exploratory and interactive mining.

• Distributed techniques for security, privacy preserving data mining Peer-to-Peer Data Mining.

• High performance data stream mining and management.

• Resource and location-aware mining algorithms.

• Data mining in mobile environments.

• Theoretical foundations for resource-aware mining in a mobile, streaming and/or distributed environment.

• Systems support for resource and location aware data mining.

• Efficient, scalable, disk-based, parallel and distributed algorithms for large-scale data mining and pre-processing and postprocessing tasks.

• Parallel or distributed frameworks for stream management, KDD systems, and parallel or distributed mining.

• Applications of parallel and distributed datamining (PDDM) in business, science, engineering, medicine, and other disciplines.

Workshop History

This is the 8th workshop on this theme held annually. Traditionally, the workshop has been held along-side the SIAM Data Mining (SDM) Conference, even if the first four editions were organized in conjunction with IPDPS, and were held at Orlando (HPDM’98), San Juan (HPDM’99), Cancun and San Francisco (PDDM’01). Over the last three years the workshop has had invited papers in the areas of mobile and locationaware data mining issues (HPDM:RLM’02), pervasive and stream datamining (HPDM:PDS’03), and grid data mining (HPDM:GRID’04).

Venue

The 2005 SIAM International Conference on Data Mining and its satellite Workshops will be held at the Sutton Place Hotel in Newport Beach, CA. The Sutton Place hotel is located in picturesque Newport Beach known as the colorful coast and conveniently situated in the heart of the Orange County business districts of Irvine and Tustin.

Workshop Co-Chairs

Salvatore Orlando, Universita Ca’ Foscari di Venezia K. Sivakumar, Washington State University

Program Committee

Gagan Agrawal, Ohio State University Rong Chen, University of Pennsylvania Chris Giannella, University of Maryland, Baltimore County Sara Graves, The University of Alabama at Huntsville Matthias Klusch, German Research Center for Art. Intell. Shonali Krishnaswamy, Monash University Michael May, Fraunhofer Institute for Autonomous Intelligent Systems Hoony Park, Oak Ridge National Laboratory Raffaele Perego, Consiglio Nazionale delle Ricerche Assaf Schuster, Technion Parthasarathy Srinivasan, Ohio State University Ashok Srivastava, NASA Ames Research Center Vaidy Sunderam, Emory University Domenico Talia, Universita della Calabria Ran Wolff, Technion Mohammed Zaki, Rensselaer Polytechnic Institute

Steering Committee

Hillol Kargupta, University of Maryland, Baltimore County Vipin Kumar, University of Minnesota Srinivasan Parthasarathy, Ohio State University David Skillicorn, Queens University Mohammed Zaki, Rensselaer Polytechnic Institute

Submission of Papers

We invite papers treating the above topics in one of many ways. The papers could describe new results, give overview or experiences with existing systems, describe new and emerging applications, present work in progress where interesting insights have been gained, or critically survey existing work. The papers should not exceed 3000 words. You can submit by emailing the PS or PDF file to

siva@eecs.wsu.edu.

Important dates:

January 7, 2005  Paper Submissions
February 4, 2005 Notification to authors
February 25, 2005 Final paper

HPDM’05: http://www.eecs.wsu.edu/∼siva/hpdm2005/
SDM’05:  http://www.siam.org/meetings/sdm05/